### v3

@@ 4.2 Competing Models
The prior distributions for the common parameters are specified as:
@@
$\nu\sim\mathcal{G}(5,0.5),\quad\beta\sim\mathcal{N}(0,1),\quad\frac{\delta+1}{2}\sim\mathcal{B}(1,1),\quad\gamma\sim G(1,1).$ (65)
@@
To further benchmark the RSV models, we also consider GARCH-type models, which are widely used to capture time-varying volatility, including ARCH (Engle, 1982), GARCH (Bollerslev, 1986), and their variants. Among these, we employ the EGARCH model introduced by Nelson (1991). Furthermore, we include the REGARCH model, a generalized version of the RGARCH framework that incorporates RV measures. See Hansen et al. (2012); Hansen and Huang (2016) for foundational work, and Takahashi et al. (2023) for an overview.
Both EGARCH and REGARCH models are estimated via the quasi-maximum likelihood (QML) method under standard normal and Student’s $t$ innovations. One-day-ahead forecasts of volatility, VaR, and ES are computed using the estimated parameters. For Bayesian approaches to RGARCH-type models and their applications to volatility and tail risk forecasting, see Chen et al. (2023).
@@ 4.3 Out-of-sample Forecasting
At each forecast point, we draw 15,000 predictive samples from the posterior predictive distribution. For the SV and RSV models, we compute both the posterior means and medians of the volatility forecasts. [^6] We report the median rather than the mean for volatility forecasts of SV models, as their posterior distributions occasionally exhibit heavy tails, which can disproportionately inflate the means. Additionally, we calculate predictive quantiles of the return distribution to obtain VaR and ES estimates. This procedure yields 603 forecasts for the DJIA and 590 for the N225, covering the period from early May 2017 to late September 2019.
@@ 4.3.1 Volatility Forecasts
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: Log-scale volatility forecasts together with the adjusted RV5 for DJIA: Overall period (top) and sub-periods (bottom).
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 7: Log-scale volatility forecasts together with the adjusted RV5 for N225: Overall period (top) and sub-periods (bottom).
Figures 6 and 7 display the log-scale volatility forecasts alongside the adjusted RV5 series for the DJIA and N225. For the DJIA, SV model forecasts—especially those from the SV-GH-ST specification—exhibit pronounced volatility spikes. Across both indices, forecast patterns are more homogeneous within model families (e.g., SV, RSV, EGARCH, REGARCH) than across different distributional assumptions, indicating that model structure plays a more dominant role than distributional form.
Table 3: QLIKE scores of volatility forecasts for DJIA and N225
@@
RSV-N $0.245^{*}$ $0.452^{*}$ $0.486^{*}$ $0.567^{*}$ $0.043^{*}$ $0.047^{*}$ $0.047^{*}$ $0.043^{*}$ $0.241$ $6.0$
RSV-T $0.231^{*}$ $0.435^{*}$ $0.471^{*}$ $0.551^{*}$ $0.043^{*}$ $0.048$ $0.047^{*}$ $0.041^{*}$ $0.234$ $4.6$
RSV-AZ-SN $0.240^{*}$ $0.449^{*}$ $0.485^{*}$ $0.566^{*}$ $0.043^{*}$ $0.048^{*}$ $0.048^{*}$ $0.043^{*}$ $0.240$ $6.1$
RSV-AZ-ST $0.227^{*}$ $0.425^{*}$ $0.462^{*}$ $0.540^{*}$ $0.043^{*}$ $0.049$ $0.048^{*}$ $0.041^{*}$ $0.229$ $4.1$
RSV-FS-SN $0.252^{*}$ $0.471$ $0.505^{*}$ $0.589$ $0.042^{*}$ $0.046^{*}$ $0.047^{*}$ $0.045^{*}$ $0.250$ $7.2$
RSV-FS-ST $0.242^{*}$ $0.452^{*}$ $0.487^{*}$ $0.569^{*}$ $0.042^{*}$ $0.047^{*}$ $0.047^{*}$ $0.043^{*}$ $0.241$ $6.0$
RSV-GH-ST $0.223^{*}$ $0.467^{*}$ $0.498^{*}$ $0.577^{*}$ $0.044$ $0.047^{*}$ $0.046^{*}$ $0.040^{*}$ $0.243$ $5.9$
EGARCH-N $0.327$ $0.441^{*}$ $0.510^{*}$ $0.575^{*}$ $0.167$ $0.175$ $0.162$ $0.141$ $0.312$ $11.0$
EGARCH-T $0.322$ $0.437^{*}$ $0.505^{*}$ $0.558^{*}$ $0.183$ $0.191$ $0.176$ $0.155$ $0.316$ $12.0$
REGARCH-N $0.248^{*}$ $0.428^{*}$ $0.476^{*}$ $0.556^{*}$ $0.058$ $0.063$ $0.060$ $0.046^{*}$ $0.242$ $7.6$
REGARCH-T $0.245^{*}$ $0.424^{*}$ $0.472^{*}$ $0.551^{*}$ $0.056$ $0.062$ $0.059$ $0.045^{*}$ $0.239$ $6.4$
@@
Table 3 reports QLIKE scores across all models and volatility proxies. An asterisk () denotes inclusion in the 90% model confidence set (MCS) of Hansen et al. (2011). [^7] The rolling window scheme satisfies the stationarity assumption required by the MCS bootstrap. We use 1,000 bootstrap replications with a block size of 10. See Hansen et al. (2011), Section 4.3. Overall, RSV and REGARCH models attain lower QLIKE scores than SV and EGARCH counterparts. For the DJIA, SV-FS-ST achieves the lowest scores under RK, BV, and Med. For the N225, RSV models dominate REGARCH, whereas the two perform comparably for the DJIA. On average, the RSV-AZ-ST model delivers the best performance across indices.
RV5 RK
Refer to caption Refer to caption
BV Med
Refer to caption Refer to caption
@@
Figure 8: Cumulative QLIKE loss differences in comparison to the SV-N model for DJIA.
RV5 RK
Refer to caption Refer to caption
BV Med
Refer to caption Refer to caption
@@
Figure 9: Cumulative QLIKE loss differences in comparison to the SV-N model for N225.
@@
To assess time-varying performance, Figures 8 and 9 depict cumulative loss differences (CLDs) relative to SV-N. Positive values indicate superior performance. When RV5 is used as the volatility proxy, the RSV and REGARCH models consistently outperform the SV and EGARCH models across the entire forecast horizon for both the DJIA and N225, underscoring the value of incorporating RV in prediction. For the DJIA, however, when alternative proxies such as RK, BV, or Med are used, the SV-FS-SN and SV-FS-ST models exhibit improved performance after mid-2018, occasionally surpassing RSV and REGARCH. In contrast, for the N225, RSV and REGARCH models maintain superior performance across all proxies, with clearer distinctions emerging between the two model classes in late 2017.
Table 4: Summary of GW tests on QLIKE of volatility forecasts for DJIA and N225
DJIA
SV-N SV-T AZ-SN AZ-ST FS-SN FS-ST GH-ST RSV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST EG-N EG-T REG-N REG-T
SV-N $90$ $12$ $77$ $98$ $96$ $58$ $94$ $96$ $94$ $96$ $92$ $94$ $97$ $67$ $71$ $92$ $93$
SV-T $-2.3$ $4$ $9$ $92$ $98$ $14$ $93$ $96$ $93$ $96$ $90$ $94$ $96$ $50$ $56$ $93$ $94$
SV-AZ-SN $2.2$ $4.3$ $94$ $96$ $97$ $79$ $97$ $98$ $97$ $98$ $95$ $97$ $98$ $71$ $74$ $95$ $96$
SV-AZ-ST $-1.4$ $1.7$ $-4.7$ $92$ $94$ $26$ $94$ $96$ $94$ $96$ $92$ $94$ $97$ $56$ $63$ $94$ $94$
SV-FS-SN $-7.3$ $-5.4$ $-7.4$ $-5.8$ $84$ $2$ $65$ $75$ $69$ $78$ $57$ $66$ $80$ $10$ $13$ $68$ $69$
SV-FS-ST $-6.2$ $-6.1$ $-6.7$ $-6.0$ $-1.1$ $2$ $60$ $72$ $65$ $77$ $54$ $61$ $80$ $7$ $9$ $64$ $67$
SV-GH-ST $-0.2$ $1.1$ $-1.1$ $0.6$ $4.5$ $4.7$ $98$ $98$ $98$ $98$ $97$ $98$ $99$ $66$ $67$ $95$ $95$
RSV-N $-5.3$ $-4.6$ $-6.0$ $-5.1$ $-1.7$ $-1.3$ $-4.7$ $100$ $97$ $99$ 00 $99$ $99$ $15$ $20$ $59$ $63$
RSV-T $-6.0$ $-5.4$ $-6.7$ $-6.0$ $-2.5$ $-2.1$ $-5.4$ $-5.6$ $3$ $81$ $3$ $1$ $91$ $9$ $14$ $39$ $44$
RSV-AZ-SN $-5.4$ $-4.8$ $-6.1$ $-5.3$ $-1.9$ $-1.5$ $-4.9$ $-1.7$ $6.6$ $98$ $9$ $25$ $97$ $13$ $18$ $52$ $57$
RSV-AZ-ST $-6.0$ $-5.6$ $-6.7$ $-6.1$ $-2.7$ $-2.4$ $-5.4$ $-5.0$ $-2.5$ $-9.6$ $6$ $8$ $91$ $9$ $13$ $29$ $35$
RSV-FS-SN $-4.9$ $-4.1$ $-5.7$ $-4.7$ $-1.2$ $-0.8$ $-4.3$ $2.7$ $8.2$ $4.3$ $6.4$ $97$ $95$ $20$ $26$ $66$ $71$
RSV-FS-ST $-5.4$ $-4.7$ $-6.2$ $-5.3$ $-1.8$ $-1.4$ $-4.8$ $-1.1$ $8.5$ $1.5$ $5.8$ $-6.6$ $94$ $15$ $18$ $56$ $59$
RSV-GH-ST $-6.2$ $-5.8$ $-7.0$ $-6.4$ $-3.0$ $-2.6$ $-5.7$ $-6.2$ $-5.0$ $-9.9$ $-4.7$ $-7.4$ $-7.2$ $8$ $11$ $24$ $28$
EG-N $-1.2$ $-0.1$ $-1.8$ $-0.6$ $4.3$ $5.4$ $-0.9$ $3.8$ $4.5$ $4.0$ $4.7$ $3.4$ $3.9$ $4.9$ $68$ $89$ $92$
EG-T $-1.3$ $-0.4$ $-1.9$ $-0.8$ $3.1$ $4.2$ $-1.0$ $3.1$ $3.8$ $3.4$ $4.0$ $2.8$ $3.2$ $4.2$ $-0.7$ $86$ $89$
REG-N $-4.6$ $-4.3$ $-5.1$ $-4.4$ $-1.4$ $-1.1$ $-4.2$ $0.2$ $1.3$ $0.6$ $1.8$ $-0.3$ $0.4$ $2.1$ $-4.1$ $-3.6$ $92$
REG-T $-4.6$ $-4.3$ $-5.1$ $-4.4$ $-1.5$ $-1.2$ $-4.2$ $0.0$ $1.1$ $0.4$ $1.5$ $-0.5$ $0.2$ $1.7$ $-4.2$ $-3.7$ $-1.9$
N225
SV-N SV-T AZ-SN AZ-ST FS-SN FS-ST GH-ST RSV-N T AZ-SN AZ-ST FS-SN FS-ST GH-ST EG-N EG-T REG-N REG-T
SV-N $4$ 00 $3$ $56$ $50$ 00 $100$ $100$ $100$ $100$ $100$ $100$ $100$ $58$ $28$ $100$ $100$
SV-T $1.3$ $83$ $66$ $99$ $88$ $81$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $89$ $56$ $100$ $100$
SV-AZ-SN $2.2$ $-0.8$ $15$ $72$ $56$ $17$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $70$ $40$ $100$ $100$
SV-AZ-ST $1.5$ $-0.1$ $1.0$ $96$ $77$ $69$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $85$ $52$ $100$ $100$
SV-FS-SN $-1.4$ $-1.7$ $-1.8$ $-1.8$ $4$ $4$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $33$ $9$ $100$ $100$
SV-FS-ST $-0.6$ $-1.5$ $-1.2$ $-1.6$ $1.4$ $1$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $74$ $13$ $100$ $100$
SV-GH-ST $1.9$ $-0.1$ $0.8$ $-0.1$ $3.0$ $2.1$ $100$ $100$ $100$ $100$ $100$ $100$ $100$ $76$ $62$ $100$ $100$
RSV-N $-5.9$ $-4.8$ $-5.7$ $-5.1$ $-6.8$ $-6.2$ $-6.7$ $1$ 00 $1$ $100$ $92$ $3$ 00 00 00 00
RSV-T $-5.8$ $-4.8$ $-5.6$ $-5.0$ $-6.8$ $-6.1$ $-6.6$ $2.5$ $100$ $2$ $100$ $100$ 00 00 00 00 00
RSV-AZ-SN $-5.9$ $-4.8$ $-5.7$ $-5.1$ $-6.8$ $-6.1$ $-6.7$ $2.8$ $-1.6$ 00 $99$ $99$ $1$ 00 00 00 00
RSV-AZ-ST $-5.8$ $-4.8$ $-5.6$ $-5.0$ $-6.8$ $-6.1$ $-6.6$ $2.3$ $0.5$ $1.5$ $100$ $100$ 00 00 00 00 00
RSV-FS-SN $-5.9$ $-4.8$ $-5.7$ $-5.1$ $-6.8$ $-6.1$ $-6.7$ $-1.9$ $-2.4$ $-2.6$ $-2.3$ 00 $2$ 00 00 00 00
RSV-FS-ST $-5.9$ $-4.8$ $-5.7$ $-5.1$ $-6.8$ $-6.1$ $-6.7$ $-0.8$ $-3.0$ $-1.9$ $-2.8$ $1.3$ $1$ 00 00 00 00
RSV-GH-ST $-5.8$ $-4.7$ $-5.6$ $-5.0$ $-6.8$ $-6.1$ $-6.6$ $4.6$ $5.8$ $4.1$ $4.9$ $3.5$ $4.7$ 00 00 00 00
EG-N $-0.7$ $-1.3$ $-1.1$ $-1.3$ $0.3$ $-0.6$ $-2.0$ $7.4$ $7.3$ $7.4$ $7.3$ $7.4$ $7.4$ $7.3$ $1$ $100$ $100$
EG-T $0.5$ $-0.5$ $0.0$ $-0.5$ $2.3$ $2.0$ $-0.5$ $6.4$ $6.4$ $6.4$ $6.4$ $6.4$ $6.4$ $6.4$ $2.7$ $100$ $100$
REG-N $-5.4$ $-4.4$ $-5.2$ $-4.7$ $-6.4$ $-5.7$ $-6.2$ $9.1$ $9.2$ $9.0$ $9.3$ $8.2$ $8.6$ $8.9$ $-6.9$ $-6.0$ $100$
REG-T $-5.5$ $-4.5$ $-5.3$ $-4.7$ $-6.4$ $-5.8$ $-6.3$ $9.0$ $9.0$ $8.8$ $9.1$ $8.0$ $8.5$ $8.7$ $-7.0$ $-6.1$ $-8.4$
  • Notes: EG and REG refer to the EGARCH and REGARCH models, respectively. The lower triangular sections report the test statistics from the unconditional GW test; a negative (positive) value indicates that the row (column) model performs better than the column (row) model. Conversely, the upper triangular sections present the percentage of predicted losses, as defined in equation (62); a lower (higher) percentage suggests that the row (column) model outperforms the column (row) model. Dark gray and light gray shading indicate rejection of the GW test at the 1% and 5% significance levels, respectively.
Table 4 presents results from the GW tests. The lower triangular part reports unconditional test statistics: negative values favor the row model. The upper triangular part shows conditional win percentages, where larger values indicate better performance by the column model. Shaded cells denote significance at the 1% (dark gray) and 5% (light gray) levels.
@@ 4.3.2 VaR and ES forecasts
Figure 10 illustrates the 1% VaR and ES forecasts for the DJIA and N225 indices, as generated by the RSV-FS-ST and REGARCH-T models. VaR violations—instances where realized returns fall below the forecasted VaR—are observed across both models, occurring during periods of market stress such as early 2018 and late 2019, as well as during more tranquil phases. For clarity, the forecasts from other models, which follow similar patterns with minor differences in scale, are omitted.
RSV-FS-ST for DJIA RSV-FS-ST for N225
@@
Figure 10: 1% VaR and ES forecasts of RSV-FS-ST and REGARCH-T models for DJIA and N225.
Table 5 presents the empirical violation rates ($\hat{\alpha}$), the average FZ0 loss values (as defined in Equation (57) of Patton et al. (2019)), and the corresponding $p$-values from the dynamic quantile (DQ) test and the MCS procedure. At the 1% level, all models tend to underestimate tail risk, as indicated by violation rates exceeding the nominal level. In contrast, for $\alpha=5\%$, the violation rates are generally well aligned with the target.
@@
The DQ test by Engle and Manganelli (2004), which regresses VaR violations on a constant, the forecasted VaR, and a one-period lagged violation indicator, suggests that the SV-FS-ST and SV-GH-ST models fail the independence test at the 1% level. This implies potential misspecification in their dynamic structures.
Table 5: Violation rates and FZ0 losses of VaR and ES forecasts for DJIA
@@
EGARCH-N $2.65$ $0.22$ $1.2928$ $0.40$ $4.81$ $0.90$ $0.6226$ $0.27$
EGARCH-T $1.82$ $0.83$ $1.0797$ $1.00$ $4.98$ $0.92$ $0.5855$ $1.00$
REGARCH-N $3.15$ $0.24$ $1.3273$ $0.25$ $5.47$ $0.81$ $0.5652$ $1.00$
REGARCH-T $2.99$ $0.23$ $1.3202$ $0.27$ $5.47$ $0.81$ $0.5621$ $1.00$
@@
Table 6 summarizes the results for the N225 index. Across both target levels, most models yield violation rates consistent with the nominal levels. The RSV-AZ-ST model again performs best in terms of FZ0 loss, and both RSV and REGARCH models are included in the 75% MCS. Notably, the SV-GH-ST and EGARCH-N models fall outside the MCS at the 5% level, indicating weaker predictive accuracy. Unlike the DJIA case, none of the models are rejected by the DQ test for N225.
Table 6: Violation rates and FZ0 losses of VaR and ES forecasts for N225
@@
EGARCH-N $1.02$ $0.93$ $1.2818$ $0.58$ $3.73$ $0.79$ $0.8450$ $0.00$
EGARCH-T $1.02$ $0.91$ $1.2344$ $0.29$ $3.73$ $0.77$ $0.8292$ $0.21$
REGARCH-N $1.69$ $0.71$ $1.1844$ $1.00$ $4.41$ $0.94$ $0.7997$ $1.00$
REGARCH-T $1.02$ $0.79$ $1.1306$ $1.00$ $4.92$ $0.99$ $0.7959$ $1.00$
@@
Figure 11 shows the CLDs relative to the SV-N benchmark. For the DJIA, REGARCH models perform poorly during the early 2018 volatility spike at the 1% level. From mid-2018 onwards, REGARCH and RSV models outperform SV and EGARCH models at the 5% level. For the N225, the superiority of RSV and REGARCH models is consistent throughout the forecast horizon, reinforcing the benefit of incorporating RV.
$\alpha=1\%$ for DJIA $\alpha=5\%$ for DJIA
Refer to caption Refer to caption
$\alpha=1\%$ for N225 $\alpha=5\%$ for N225
Refer to caption Refer to caption
Figure 11: Cumulative loss differences (FZ0) relative to the SV-N model for DJIA and N225.
@@
Table 7 reports the GW test results for the FZ0 loss function for the DJIA. The interpretation of the lower and upper triangular elements follows that of Table 4. The main findings are summarized below.
@@
At the 1% level, the unconditional GW test reveals that SV-AZ-SN significantly outperforms SV-N, while SV-GH-ST is significantly outperformed by SV-AZ-SN. Among RSV models, RSV-T, RSV-AZ-SN, RSV-AZ-ST, RSV-FS-SN, RSV-FS-ST, and RSV-GH-ST all show significantly better predictive accuracy than RSV-N. For EGARCH (EG) models, EG-T significantly outperforms EG-N. Regarding REGARCH (REG) models, both REG-N and REG-T are significantly outperformed by several RSV models, particularly RSV-AZ-SN, RSV-AZ-ST, RSV-FS-ST, and RSV-GH-ST.
@@
The conditional GW test at the 1% level indicates that SV-T performs significantly worse than SV-AZ-SN, but better than EG-N. RSV-N is significantly outperformed by RSV-T, RSV-AZ-SN, RSV-FS-SN, and RSV-FS-ST. Moreover, all RSV variants (RSV-AZ-SN, RSV-AZ-ST, RSV-FS-SN, RSV-FS-ST, RSV-GH-ST) significantly outperform REG-N and REG-T.

### v4

@@ 4.2 Competing Models
The prior distributions for the common parameters are specified as follows:
@@
These priors are chosen to be weakly informative and follow standard specifications commonly adopted in the stochastic volatility literature, allowing the data to dominate posterior inference.
@@
$\nu\sim\mathcal{G}(5,0.5)\mathbbm{1}\{\nu>4\},\quad\beta\sim\mathcal{N}(0,1),\quad\frac{\delta+1}{2}\sim\mathcal{B}(1,1),\quad\gamma\sim G(1,1).$ (76)
The truncation $\nu>4$ ensures the existence of the fourth moment, which is required for volatility forecasting and risk evaluation.
@@
To further benchmark the RSV models, we also consider the EGARCH and REGARCH models described in Sections 2.8 and 2.9, respectively. These models are estimated within a Bayesian framework using the MH algorithm under standard normal and Student’s $t$ innovations. One-day-ahead forecasts of volatility, VaR, and ES are then obtained based on the posterior draws. For other Bayesian approaches to RGARCH-type models and their applications to volatility and tail risk forecasting, see Chen et al. (2023).
@@ 4.3 Out-of-sample Forecasting
At each forecast point, we draw 15,000 predictive samples from the posterior predictive distribution. For the SV and RSV models, we compute both the posterior means and medians of the volatility forecasts. [^8] We report the median rather than the mean for volatility forecasts of SV models, as the posterior predictive distribution of the one-step-ahead volatility forecast occasionally exhibits heavy tails, which can disproportionately inflate the posterior mean. Additionally, we calculate predictive quantiles of the return distribution to obtain VaR and ES estimates. This procedure yields 603 forecasts for the DJIA and 590 for the N225, covering the period from early May 2017 to late September 2019.
@@ 4.3.1 Volatility Forecasts
Figure 6: Log-scale volatility forecasts together with the adjusted RV5 for the DJIA.
Figure 7: Log-scale volatility forecasts together with the adjusted RV5 for the N225.
Figures 6 and 7 display the log-scale volatility forecasts alongside the adjusted RV5 series for the DJIA and N225. For the DJIA, SV model forecasts—especially those from the SV-GH-ST specification—exhibit pronounced volatility spikes. Across both indices, forecast patterns are more homogeneous within model families (e.g., SV, RSV, EGARCH, REGARCH) than across different distributional assumptions, indicating that model structure plays a more dominant role than distributional form. Figures with alternative realized volatility proxies are omitted for brevity, as the volatility forecasts are identical.
Table 3: QLIKE scores of volatility forecasts for the DJIA and N225.
@@
RSV-N $0.245^{*}$ $0.454$ $0.489^{*}$ $0.571$ $043^{*} 0.0$470.047 $047^{*} 0.0$$043^{*}$ $0.242$ $7.1$
RSV-T $0.231^{*}$ $0.433^{*}$ $0.469^{*}$ $548^{*} 0.0$$43^{*} 0.0$480.048 $047^{*} 0.0$$041^{*}$ $0.233$ $5.9$
RSV-AZ-SN $0.240^{*}$ $0.449$ $0.485^{*}$ $0.566$ $043^{*} 0.0$480.048 $047^{*} 0.0$$043^{*}$ $0.240$ $6.2$
RSV-AZ-ST $0.227^{*}$ $0.425^{*}$ $0.462^{*}$ $540^{*} 0.0$$43^{*} 0.0$490.049 $048^{*} 0.0$$041^{*}$ $0.229$ $5.0$
RSV-FS-SN $0.252^{*}$ $0.471$ $0.505$ $0.589$ $042^{*} 0.0$$46^{*} 0.0$$47^{*} 0.0$450.045 $0.250$ $7.5$
RSV-FS-ST $0.242^{*}$ $0.452$ $0.487^{*}$ $0.569$ $042^{*} 0.0$470.047 $047^{*} 0.0$$043^{*}$ $0.241$ $6.4$
RSV-GH-ST $0.223^{*}$ $0.416^{*}$ $0.453^{*}$ $531^{*} 0.0$440.044 $0.049$ $0.048$ $0.041^{*}$ $0.226$ $4.9$
EGARCH-N $0.321$ $0.431^{*}$ $0.499^{*}$ $0.563^{*}$ $0.166$ $0.174$ $0.161$ $0.141$ $0.307$ $10.6$
EGARCH-T $0.303^{*}$ $0.401^{*}$ $0.465^{*}$ $0.516^{*}$ $0.184$ $0.193$ $0.178$ $0.155$ $0.299$ $11.0$
REGARCH-N $0.240^{*}$ $0.417^{*}$ $0.464^{*}$ $543^{*} 0.0$570.057 $0.063$ $0.060$ $0.045^{*}$ $0.236$ $7.2$
REGARCH-T $0.221^{*}$ $0.382^{*}$ $0.428^{*}$ $504^{*} 0.0$570.057 $0.063$ $0.060$ $0.044^{*}$ $0.220$ $5.0$
@@
Table 3 reports QLIKE scores across all models and volatility proxies. An asterisk () denotes inclusion in the 90% model confidence set (MCS) of Hansen et al. (2011). [^9] The rolling window scheme satisfies the stationarity assumption required by the MCS bootstrap. We use 1,000 bootstrap replications with a block size of 10. See Hansen et al. (2011), Section 4.3. Overall, RSV and REGARCH models attain lower QLIKE scores than SV and EGARCH counterparts. For the DJIA, REGARCH-T achieves the lowest scores under RV5 and BV, while SV-FS-ST achieves the lowest scores under RK and Med. For the N225, RSV models dominate REGARCH models. On average, the RSV-GH-ST model delivers the best performance across indices, followed by the RSV-AZ-ST and REGARCH-T models, with very small margins.
@@
Figure 8: Cumulative QLIKE loss differences relative to the SV-N model, based on RV5 as the volatility proxy, for the DJIA.
@@
Figure 9: Cumulative QLIKE loss differences relative to the SV-N model, based on RV5 as the volatility proxy, for the N225.
@@
To assess time-varying performance, Figures 8 and 9 depict cumulative loss differences (CLDs) relative to SV-N, based on RV5 as the volatility proxy. [^10] Results based on alternative realized volatility proxies are reported in the Supplementary Material. Positive values indicate superior performance. The RSV and REGARCH models consistently outperform the SV and EGARCH models across the entire forecast horizon for both the DJIA and N225, underscoring the value of incorporating realized volatility in prediction. In particular, the performance advantage of RSV and REGARCH is stable over time, with clear and persistent separations from the SV and EGARCH benchmarks in both markets.
Figure 10: Pairwise comparison of GW tests on QLIKE of volatility forecasts, based on RV5 as the volatility proxy, for the DJIA. The lower triangular part reports unconditional GW test statistics, while the upper triangular part shows win proportions as defined in Equation (72). Lower values (green) indicate superior performance of the row model, whereas higher values (orange) favor the column model. Double and single asterisks (∗∗ and ) denote statistical significance at the 1% and 5% levels, respectively, based on unconditional GW test p-values for the lower triangular part and conditional GW test p-values for the upper triangular part.
Figures 10 and 11 visualize the GW test results, based on RV5 as the volatility proxy, using heatmaps. [^11] Results based on alternative realized volatility proxies are reported in the Supplementary Material. The lower triangular part presents unconditional GW test statistics, whereas the upper triangular part reports win proportions as defined in Equation (72). Lower values (green) favor the row model, while larger values (orange) indicate better performance by the column model. Double and single asterisks (** and *) denote significance at the 1% and 5% levels, respectively, based on unconditional and conditional GW test p-values for the lower and upper triangular parts.
@@
Figure 11: Pairwise comparison of GW tests on QLIKE of volatility forecasts, based on RV5 as the volatility proxy, for the N225. See Figure 10 for additional details.
@@ 4.3.2 VaR and ES forecasts
Figure 12 illustrates the 1% VaR and ES forecasts for the DJIA and N225 indices, as generated by the RSV-AZ-ST and REGARCH-T models. VaR violations—instances where realized returns fall below the forecasted VaR—are observed across both models, occurring during periods of market stress such as early 2018 and late 2019, as well as during more tranquil phases. For clarity, the forecasts from other models, which follow similar patterns with minor differences in scale, are omitted.
RSV-AZ-ST for DJIA RSV-AZ-ST for N225
@@
Figure 12: 1% VaR and ES forecasts of RSV-FS-ST and REGARCH-T models for the DJIA and N225.
Table 4 presents the empirical violation rates ($\hat{\alpha}$), the average FZ0 loss values (as defined in Equation (67)), and the corresponding $p$-values from the dynamic quantile (DQ) test of Engle and Manganelli (2004) and the MCS procedure for the DJIA. At the 1% level, all models tend to underestimate tail risk, as indicated by violation rates exceeding the nominal level. In contrast, for $\alpha=5\%$, the violation rates are generally well aligned with the target.
@@
The DQ test, which regresses VaR violations on a constant, the forecasted VaR, and a one-period lagged violation indicator, suggests that the SV-FS-ST and SV-GH-ST models fail the independence test at the 1% level. This implies potential misspecification in their dynamic structures.
Table 4: Violation rates and FZ0 losses of VaR and ES forecasts for the DJIA.
@@
EGARCH-N $2.49$ $0.21$ $1.2822$ $0.35$ $4.64$ $0.85$ $0.6206$ $0.27$
EGARCH-T $1.49$ $0.92$ $1.0605$ $1.00$ $4.98$ $0.91$ $0.5826$ $1.00$
REGARCH-N $3.15$ $0.24$ $1.3048$ $0.31$ $5.31$ $0.78$ $0.5580$ $1.00$
REGARCH-T $2.16$ $0.16$ $1.0476$ $1.00$ $5.47$ $0.81$ $0.5413$ $1.00$
@@
Table 5 summarizes the results for the N225. Across both target levels, most models yield violation rates consistent with the nominal levels. The RSV-AZ-ST model again performs best in terms of FZ0 loss, and both RSV and REGARCH models are included in the 75% MCS. Notably, the SV-GH-ST and EGARCH-N models fall outside the MCS at the 5% level, while the EGARCH-T model remains outside the MCS even at the 1% level, indicating weaker predictive accuracy. Unlike the DJIA case, none of the models are rejected by the DQ test for the N225.
Table 5: Violation rates and FZ0 losses of VaR and ES forecasts for the N225.
@@
EGARCH-N $1.02$ $0.95$ $1.2944$ $0.49$ $3.56$ $0.72$ $0.8480$ $0.00$
EGARCH-T $1.02$ $0.90$ $1.2448$ $0.17$ $3.56$ $0.67$ $0.8319$ $0.15$
REGARCH-N $1.69$ $0.70$ $1.1744$ $1.00$ $4.41$ $0.94$ $0.7988$ $1.00$
REGARCH-T $1.02$ $0.78$ $1.1343$ $1.00$ $4.75$ $0.98$ $0.7922$ $1.00$
@@
Figures 13 and 14 show the CLDs relative to the SV-N benchmark for the DJIA. At the 1% level, REGARCH models perform poorly during the early 2018 volatility spike. From mid-2018 onwards, however, REGARCH and RSV models consistently outperform SV and EGARCH models at the 5% level, highlighting the benefits of incorporating realized volatility into tail risk forecasting. Consistent results are obtained for the N225, reinforcing the benefit of incorporating RV, and are therefore reported in the Supplementary Material.
Figure 13: Cumulative loss differences (FZ0) relative to the SV-N model ($\alpha=1\%$) for the DJIA.
Figure 14: Cumulative loss differences (FZ0) relative to the SV-N model ($\alpha=5\%$) for the DJIA.
@@
Figure 15: Pairwise comparison of GW tests on FZ0 loss of VaR and ES forecasts ($\alpha=1\%$) for the DJIA. See Figure 10 for additional details.
Figure 16: Pairwise comparison of GW tests on FZ0 loss of VaR and ES forecasts ($\alpha=5\%$) for the DJIA. See Figure 10 for additional details.
@@
Figures 15 and 16 visualize the GW test results based on the FZ0 loss function for the DJIA using heatmaps. The interpretation of the lower and upper triangular elements follows that of volatility forecasts in Section 4.3.1. The main findings are summarized below.
@@
At the 1% level, the unconditional GW test reveals that SV-AZ-SN significantly outperforms SV-N, while SV-GH-ST is significantly outperformed by SV-AZ-SN. Among RSV models, RSV-N exhibits inferior predictive accuracy relative to all other RSV specifications. For EGARCH models, EGARCH-T significantly outperforms EGARCH-N, while REGARCH models are significantly outperformed by several RSV specifications.
@@
The conditional GW test at the 1% level indicates that SV-T performs significantly worse than SV-AZ-SN, but better than EGARCH-N. RSV-N is significantly outperformed by RSV-T, RSV-AZ-SN, RSV-FS-SN, and RSV-FS-ST. Moreover, RSV models with skew(-t) distributions significantly outperform REGARCH-N and REGARCH-T.